OpenStudy (anonymous):

square root 98 divided by square root 2

OpenStudy (anonymous):

\[\frac{\sqrt{98}}{\sqrt{2}} = \sqrt{\frac{98}{2}}\]

OpenStudy (anonymous):

u have to divide them , thats not the anwser

OpenStudy (anonymous):

you should divide them by yourself then, I dont give you the answer directly

OpenStudy (anonymous):

i cant divide them if i dont have a calculator......

OpenStudy (anonymous):

I have already simplified the form, you should be able to divide what inside the square root now

OpenStudy (anonymous):

i know its close to 3 but i dont know if its exact

OpenStudy (anonymous):

i dont have a calculator to do it thats why i was asking the dang question.

OpenStudy (anonymous):

you don't need a calculator to do this, what is 98 : 2 ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

49....so 7

OpenStudy (anonymous):

thats it